OpenStudy (anonymous):

Write the slope-intercept equation of the line containing the points(-4,7) and (-6,-1)

OpenStudy (anonymous):

\[\frac{y-y_1}{x-x_1}=\frac{y_2-y_1}{x_2-x_1}\] x_1=-4 x_2=-6 y_1=7 y_2=-1